OpenStudy (anonymous):

Given the linear equation y = 3x + 1, find the y-coordinates of the points (1, ), (-2, ), (3, ). Please show all of your work. Plot these points and graph the linear equation. Submit your graph through the dropbox

OpenStudy (anonymous):

For the first one, substitute 1 for x and we get y = 3(1) + 1 = 4 Can you do the same for -2 and 3?

OpenStudy (anonymous):

idk

OpenStudy (earthcitizen):

Given y=mx+c for a straight line graph the points are (x,y)

OpenStudy (earthcitizen):

http://openstudy.com/#/updates/4efd0b82e4b01ad20b529f9a this might help